OpenStudy (spicytaco1):

What is the value of the expression? [(1+5)⋅2−5]⋅2

OpenStudy (viper172):

18 im guessing...

jabez177 (jabez177):

Brackets first. (1-5) x 2-5 Inside the brackets is a set of parentheses so we do the parentheses first. :) 1 - 5 = -4 We still have the 2 -5 so 2 - 5 is -3 So now our equation is [-4 x -3] x 2 Brackets first again. :) -4 x -3 = 12 because a negative x a negative ALWAYS equals a positive :) [12] x 2 = what? That will be your answer. :)

OpenStudy (retireed):

Typo it is plus not a subtract

OpenStudy (retireed):

[(1+5)⋅2−5]⋅2 6 times 2 = 12 12 - 5 = 7 then 7 times 2

jabez177 (jabez177):

Sorry!!! D: Man...

OpenStudy (retireed):

np